# Unitary discriminants of characters

### v1

@@ 6 Perfect groups with even center
    If Weierstrass-$\wp$ is ramified in $L/K$ and $d$ is the sum of the dimensions of the orthogonal constituents of $\chi\pmod{\wp}$ then Weierstrass-$\wp$ is ramified in $\Delta(\chi)$ if and only if $(-1)^{d/2}$ is not a square modulo Weierstrass-$\wp$.
Proof.
(a) For any non-dyadic prime Weierstrass-$\wp$ of $K$ the central element $z\in G$ of order $2$ acts as $-id$ on all Weierstrass-$\wp$-modular constituents of $\chi$. As $G$ is perfect, $\det(-id)=1$ and thus all these constituents have even degree. Hence $\chi\pmod{\wp}$ is unitary stable. Now (a) follows from Remark 4.7 and Theorem 5.4.
(b) As in (a) all Weierstrass-$\wp$-modular constituents of $\chi$ have even degree. Let $P$ be the prime of $L$ such that $P^{2}=\wp{\mathbb{Z}}_{L}$. By Proposition 3.10 there is a ${\mathbb{Z}}_{L_{P}}G$-lattice $\Lambda$ such that $P\Lambda^{*}\subset\Lambda\subset\Lambda^{*}$. Moreover $(\Lambda/P\Lambda^{*},\overline{H})$ is a non-degenerate symmetric bilinear space over the residue field ${\mathbb{Z}}_{L}/P\cong{\mathbb{Z}}_{K}/\wp$ and $(\Lambda^{*}/\Lambda,\tilde{H})$ is a symplectic ${\mathbb{Z}}_{L}/P{\mathbb{Z}}_{L}$-module. In particular all orthogonal Weierstrass-$\wp$-modular constituents occur with even multiplicity in $\Lambda^{*}/\Lambda$, so
$d\equiv\dim(\Lambda/P\Lambda^{*})\pmod{4}.$
Clearly $z$ acts as $-id$ on $\Lambda/P\Lambda^{*}$. Since $G$ is perfect, the Spinor norm of $-id$ is trivial. But the Spinor norm of $-id$ is the determinant of $\overline{H}$ (see for instance [10, Section 3.1.2]) and so also this determinant is trivial and $\operatorname{disc}(\overline{H})=(-1)^{d/2}\in\{1,-1\}$. Now Corollary 5.6 tells us that Weierstrass-$\wp$ is ramified in $\Delta(\chi)$, if and only if $(-1)^{d/2}$ is not a square modulo Weierstrass-$\wp$. ∎
Corollary 6.2.
Let $G$ be a perfect group $G$ whose center has an order divisible by $4$ and let $\chi\in\operatorname{Irr}^{o}(G)$ be faithful. Then only dyadic primes of $K$ might ramify in $\Delta(\chi)$.
Proof.
Let $z$ be a central element of order $4$. Then $z$ acts as a primitive fourth root of unity on all irreducible representations of $G$. Again, the determinant of such a representation is $1$, so all faithful irreducible characters have a degree that is a multiple of $4$. As this also holds for $\chi\in\operatorname{Irr}^{o}(G)$, no infinite place of $K$ ramifies in $\Delta(\chi)$. Proposition 6.1 shows that the only primes of $K$ that might possibly ramify in $\Delta(\chi)$ are the dyadic primes of $K$. ∎
6.1 The $Q_{8}$-trick
Remark 6.3.
Let $G$ be a group with cyclic center of even order and let $z\in G$ denote the central element of order 2 in $G$. Assume that $G$ contains a subgroup $U\cong Q_{8}$ with $z\in U$. Let $\chi$ be the unique faithful irreducible character of $U$. Then the restriction of any faithful character $\operatorname{{\mathfrak{X}}}$ of $G$ to $U$ is a multiple of of $\chi$ and hence unitary stable. The methods from Section 4.3 allow to read off the unitary discriminant of $\operatorname{{\mathfrak{X}}}$ just from its degree and its character field: Put $L:={\mathbb{Q}}(\operatorname{{\mathfrak{X}}})$ and let $K$ denote its maximal real subfield. Then $(-1,-1)_{K}=(L,-d)_{K}$ for $d=\operatorname{disc}_{L}((-1,-1)_{K})$ and
[…]
In both cases the discriminant algebra of $\operatorname{{\mathfrak{X}}}$ is
$\Delta(\operatorname{{\mathfrak{X}}})=[(-1,-1)_{K}]^{\operatorname{{\mathfrak{X}}}(1)/2}\in\operatorname{Br}_{2}(L,K).$
7 Antiadjoint automorphisms
In this section we assume that the characteristic of the field $K$ is not 2. For quadratic $KG$-modules $(V,Q)$ the paper [12] works out a very useful method to compute the determinant of the associated $G$-invariant symmetric bilinear form
$B:V\times V\to K,B(x,y):=Q(x+y)-Q(x)-Q(y).$
If there is an invertible $X\in\operatorname{End}_{K}(V)$ such that $X^{ad}=-X$ (adjoint with respect to $B$) then $\det(B)=\det(X)(K^{\times})^{2}$. There is an elementary proof by choosing a basis of $V$ and considering the matrices $X$ and $B$ with respect to this basis. Then
$X^{ad}=BX^{tr}B^{-1}=-X\mbox{ and hence }(XB)=-(XB)^{tr}$
is a skew symmetric matrix. As the determinant of a skew symmetric matrix is always a square we get $\det(X)=\det(B)$ modulo squares.
Note that the adjoint involution of $B$ inverts the group elements so we can search for skew adjoint invertible elements in the subalgebra of $\operatorname{End}_{K}(V)$ spanned by the group elements without having the $G$-invariant quadratic form $Q$.
For any central simple $K$-algebra $A$ there is a field extension $E$, such that $E\otimes_{K}A\cong E^{n\times n}$. For $a\in A\subseteq E\otimes_{K}A\cong E^{n\times n}$ we define the reduced norm of $a$ as the determinant of the matrix $a\in E^{n\times n}$, $N_{red}(a)=\det(a)$ (see [15, Section 9]). This definition depends neither on the chosen splitting field $E$ nor on the identification with a matrix ring. Moreover for any $a\in A$ we have $N_{red}(a)\in K$. We briefly recall the notion of a discriminant of an orthogonal involution.
Remark 7.1.
([9, Definition 7.2]) Let $K$ be a field of characteristic $\neq 2$ and $A$ be a central simple $K$-algebra of even degree $2m$, so $\dim_{K}(A)=(2m)^{2}$. An orthogonal involution $\iota$ on $A$ is a $K$-algebra anti-automorphism $\iota$ of order 2 such that the $K$-subspace
$A_{-}:=\{a\in A\mid\iota(a)=-a\}$
has dimension $m(2m-1)$. Then $A_{-}$ contains invertible elements. The discriminant of $\iota$ is defined as the square class
$\operatorname{disc}(\iota):=(-1)^{m}N_{red}(a)(K^{\times})^{2}$
for any invertible $a\in A_{-}$. In particular for $A=K^{2m\times 2m}$ and a symmetric $B\in\operatorname{GL}_{2m}(K)$ such that $\iota=\iota_{B}$ is the adjoint involution with respect to $B$, then
The adjoint involution of an $L/K$-Hermitian form $H$ is an involution of second kind on the endomorphism ring, as it restricts non trivially to the center $L$. Here the discriminant algebra from Remark 4.8 is hard to compute in general. Also it only singles out a class of $K^{\times}/N_{L/K}(L^{\times})$ which, for number fields $K$, cannot be determined by looking at reductions modulo prime ideals of $K$, as for finite fields norms are surjective. To apply similar methods as for quadratic forms to Hermitian $LG$-modules $(V,H)$, we need to find an orthogonal subalgebra of $\operatorname{End}_{L}(V)$:
Definition 7.2.
Let ${\mathcal{A}}$ be a central simple $L$-algebra of even degree $2m$ and with $L/K$-involution $\iota$ of second kind. Then a $K$-subalgebra $A$ of ${\mathcal{A}}$ is called an orthogonal subalgebra of $({\mathcal{A}},\iota)$, if and only if
  • (a)
    $A$ is a central simple $K$-algebra with $LA={\mathcal{A}}$.
  • (b)
    $A$ is invariant under $\iota$, i.e. $\iota(A)=A$.
@@ 7 Antiadjoint automorphisms
(see also [9, Proposition (10.33)]) Let $(V,H)$ be a Hermitian $L$-space of even dimension $2m$, ${\mathcal{A}}:=\operatorname{End}_{L}(V)$, and $\iota=\iota_{H}$ the adjoint involution of the non-degenerate Hermitian form $H$. Let $A$ be an orthogonal subalgebra of ${\mathcal{A}}$.
@@
    $[A]\in\operatorname{Br}_{2}(L,K)$.
@@
    $\operatorname{disc}(H)=\operatorname{disc}_{L}([A])^{m}\operatorname{disc}(\iota_{|A})\in K^{\times}/N_{L/K}(L^{\times}).$
@@
(a) Let $A$ be an orthogonal subalgebra of ${\mathcal{A}}$. Then $A$ is a central simple $K$-algebra such that $A\otimes_{K}L\cong L^{2m\times 2m}$, i.e. $A\cong{\mathcal{Q}}^{m\times m}$ for some quaternion algebra ${\mathcal{Q}}$ with center $K$ that is split by $L=K[\sqrt{-\delta}]$.
(b) Choose $b\in\operatorname{disc}_{L}([A])$. Then ${\mathcal{Q}}=(L,b)_{K}$. Consider the case $m=1$. Then
${\mathcal{Q}}=\langle i:=\operatorname{diag}(\sqrt{-\delta},-\sqrt{-\delta}),j:=\left(\begin{array}[]{cc}0&1\\ b&0\end{array}\right)\rangle_{K}\subseteq L^{2\times 2}$
with is invariant under the adjoint involution $\iota$ of the $L/K$-Hermitian form $B:=\operatorname{diag}(1,b)$. This involution is orthogonal on ${\mathcal{Q}}$ and $i$ is a skew-adjoint element so
$\operatorname{disc}(\iota_{|{\mathcal{Q}}})=-\det(i)=-\delta$
whereas $\operatorname{disc}(B)=-b$. As $\delta=N_{L/K}(\sqrt{-\delta})$ is a norm and $\operatorname{disc}(B)=\operatorname{disc}(\iota)$ we have
$\operatorname{disc}(\iota)=\operatorname{disc}_{L}([A])^{m}\operatorname{disc}(\iota_{|A}).$ (1)
For general $m$ choose a suitable $L$-basis of $V$ on which $A$ acts as block matrices with blocks in ${\mathcal{Q}}$ as above. Then the Hermitian adjoint involution $\iota$ of the block diagonal matrix ${\bf B}:=\operatorname{diag}(B,\ldots,B)$ restrict to an orthogonal involution of $A$. As $A$ is a central simple $K$-algebra, Skolem-Noether theorem tells us that any other orthogonal involution on $A$ is of the form $\kappa_{u}\iota$, where $\kappa_{u}$ is the conjugation by some symmetric unit $u\in A$. This changes both, $\operatorname{disc}(u{\bf B})=\det(u)\operatorname{disc}({\bf B})$ and $\operatorname{disc}(\iota_{|A})$ by $\det(u)$, so equation (1) holds for all Hermitian involutions $\iota=\iota_{H}$. ∎

### v2

@@ 4.2 Orthogonal subalgebras
    AA is invariant under $\iota$, i.e. $\iota(A)=A$.
@@
Note that orthogonal subalgebras do not exist in general. However for split algebras, there are always orthogonal subalgebras.
@@
([10, Proposition (10.33)], [17, Proposition 11]) Let $\mathcal{A}$} be a central simple $L$-algebra of even degree $2n$ and with $L/K$-involution $\iota$ of second kind.
@@
    Let $A$ be an orthogonal subalgebra of $({\mathcal{A}},\iota)$. Then
    $\Delta(({\mathcal{A}},\iota))]=[A]^{n}[(L,\operatorname{disc}(\iota_{|A}))]\in\operatorname{Br}(K).$
@@
    If $A$ is a symplectic subalgebra of $({\mathcal{A}},\iota)$, then
    $\Delta(({\mathcal{A}},\iota))]=[A]^{n}.$
If $({\mathcal{A}},\iota)\cong(L^{2n\times 2n},\iota_{H})$ is split, then the formulas in Theorem 4.8 applied to the discriminant $\operatorname{disc}(H)$ of the Hermitian form read as
  • (a)
    $\operatorname{disc}(H)=(-1)^{n}\operatorname{disc}_{L}([A]^{n})N_{red}(X)$ for any $X=-\iota(X)\in A^{\times}$ in the orthogonal case and
  • (b)
    $\operatorname{disc}(H)=\operatorname{disc}_{L}([A]^{n})$ in the symplectic case.
4.3 Stable subalgebras
We focus on involutions of the second kind, similar results also hold for orthogonal involutions (see [15, Theorem 5.13]). So let $L$ be a field of characteristic $\neq 2$ and let $\mathcal{A}$} be a central simple $L$-algebra of even degree and with an involution $\iota$ of the second kind.
Definition 4.9.
A semisimple subalgebra $A\leq{\mathcal{A}}$ is called stable if
  • $\iota(A)=A$ and
  • all $\iota$-invariant simple direct summands of $A$ have even degree.
So a semisimple involution invariant subalgebra
$A=\bigoplus_{i=1}^{h}A_{i}=\bigoplus_{i=1}^{h}e_{i}A$
is stable, if for all central primitive idempotents $e_{i}$ with $\iota(e_{i})=e_{i}$, the dimension $\dim_{Z(A_{i})}(A_{i})$ is even. If $\iota(e_{i})=e_{j}$ for some $j\neq i$ then the restriction of $\iota$ to $(e_{i}+e_{j})A$ is hyperbolic and has no contribution to the discriminant algebra of $({\mathcal{A}},\iota)$. The next remark shows that, up to hyperbolic involutions, we can reduce to the simple case by replacing $\mathcal{A}$} by $e_{i}{\mathcal{A}}e_{i}$.
Remark 4.10.
If ${\mathcal{A}}={\mathcal{D}}^{n\times n}$ is a central simple $L$-algebra ($\mathcal{D}$} a central $L$-division algebra) and $e^{2}=e\in{\mathcal{A}}$, then $e{\mathcal{A}}e\cong{\mathcal{D}}^{x\times x}$ for some $x\leq n$. An easy proof uses the fact that $\mathcal{A}$} has a unique simple module, ${\mathcal{D}}^{n}$ and $e{\mathcal{A}}$ is a direct sum of simple modules, so isomorphic to $\oplus^{x}{\mathcal{D}}^{n}$. Then
$(e{\mathcal{A}}e)^{op}\cong\operatorname{End}_{\mathcal{A}}(e{\mathcal{A}})\cong({\mathcal{D}}^{op})^{x\times x}$
implies that $e{\mathcal{A}}e\cong{\mathcal{D}}^{x\times x}$.
So if $e^{2}=e=\iota(e)\in A$ is an involution invariant central primitive idempotent
  • (a)
    eAeeAe is a simple algebra in $e{\mathcal{A}}e$ and
  • (b)
    $e{\mathcal{A}}e\cong{\mathcal{D}}^{x\times x}$ is a central simple $L$-algebra Brauer equivalent to $\mathcal{A}$}.
To handle reducible characters, we need to define the discriminant of a module. The idea is that the discriminant algebra of a central simple stable algebra with involution is the discriminant algebra of its unique simple module. Further this discriminant algebra behaves multiplicative with respect to direct sums of modules and hence it is enough to deal with the case that $A$ is a simple stable subalgebra of $\mathcal{A}$}.
Remark 4.11.
Assume that $A$ is a simple stable subalgebra of $\mathcal{A}$} such that $Z(A)\leq Z({\mathcal{A}})=L$. Then $LA$ is a central simple $L$-algebra with an involution $\iota_{|LA}$ of the second kind. Let $m$ be the multiplicity of the simple $LA$-module in the restriction of the simple $\mathcal{A}$}-module to $LA$. Then
$\Delta(({\mathcal{A}},\iota))]=[\Delta((LA,\iota_{|LA}))]^{m}.$
It remains to consider the situation where $E:=Z(A)>Z({\mathcal{A}})=L$. Here we assume that the characteristic of $L$ is 0, to avoid inseparable extensions. Let $F:=\operatorname{Fix}_{\iota}(E)$ denote the fixed field of $\iota$ in $E$. Then $F\geq K$ is a separable finite extension of $K$ and hence there is a well defined map, known as corestriction [23],
$N_{F/K}:\operatorname{Br}(F)\to\operatorname{Br}(K).$
In particular $N_{F/K}([(E,\delta)_{F}])=[(L,N_{F/K}(\delta))_{K}]$ by [23, Theorem 3.2] (and the fact that $E=LF=F[\ell]$ for some $\ell\in L$).
Proposition 4.12.
Let $({\mathcal{A}},\iota)$ be a central simple $L$-algebra of even degree and with $L/K$-Hermitian involution $\iota$. Let $L\leq E\leq{\mathcal{A}}$ be an extension field of $L$, such that $\iota(E)=E$. Let $F$ denote the fixed field of $\iota$ in $E$. Put $A:=C_{{\mathcal{A}}}(E)$ to denote the centraliser of $E$ in $\mathcal{A}$}. Then $(A,\iota)$ is a central simple $E$-algebra with involution. If the degree of $A$ is even, then
$\Delta(({\mathcal{A}},\iota))]=N_{F/K}([\Delta((A,\iota))]).$
@@ 4.3 Stable subalgebras
By Remark 4.5 we can assume that ${\mathcal{A}}\cong L^{2n\times 2n}$ is split. Then also $A\cong E^{2d\times 2d}$ is split, where $n/d=[E:L]$. So there is an $E/F$-Hermitian form $H\in A$ such that $\iota_{|A}=\iota_{H}$. Let $\varphi:E\hookrightarrow L^{n/d\times n/d}$ denote the regular representation. Applying $\varphi$ to the entries of the matrices yields an embedding $\phi$ of $E^{2d\times 2d}$ into $L^{2n\times 2n}$. Then $\det(\phi(H))=N_{F/K}(\det(H))$, so
$\Delta(({\mathcal{A}},\iota))]=[(L,\operatorname{disc}(\phi(H))_{K}]=[(L,N_{F/K}(\operatorname{disc}(H)))_{K}]=N_{F/K}([\Delta((A,\iota))]).$